### v1

# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representational Alignment

Abstract
Enforcing alignment between the internal representations of diffusion or flow-based generative models and those of pretrained self-supervised encoders has recently been shown to provide a powerful inductive bias, improving both convergence and sample quality. In this work, we extend this idea to inverse problems, where pretrained generative models are employed as priors. We propose applying representation alignment (Repa) between diffusion or flow-based models and a pretrained self-supervised visual encoder, such as DINOv2, to guide the reconstruction process at inference time. Although ground-truth signals are unavailable in inverse problems, we show that aligning model representations with approximate target features can substantially enhance reconstruction fidelity and perceptual realism. We provide theoretical results showing (a) the relation between the Repa regularization and a divergence measure in the DINOv2 embedding space, and (b) how Repa updates steer the model’s internal representations toward those of the clean image. These results offer insights into the role of Repa in improving perceptual fidelity. Finally, we demonstrate the generality of our approach by integrating it into multiple state-of-the-art inverse problem solvers. Extensive experiments on super-resolution, box inpainting, Gaussian deblurring, and motion deblurring confirm that our method consistently improves reconstruction quality across tasks, while also providing substantial efficiency gains reducing the number of required discretization steps without compromising the performance of the underlying solver. Our code is available at https://github.com/Sfountouris/Align_And_Invert.
1 Introduction
Pretrained diffusion and flow-based models, (sohl2015deep; ho2020denoising; esser2024scaling; lipman2023flow), have recently been at the heart of methods focusing on addressing various types of inverse problems. The crux of these approaches is to perform diffusion sampling, incorporating measurement information during the reverse process, with the goal to reconstruct the clean image at the final step, (patel2024steering; chung2023diffusion). Undoubtedly, diffusion and flow-based models have pushed the boundaries in addressing a wealth of inverse problems, (daras2024survey). However, they still struggle to produce highly detailed images particularly in cases of severe degradation and in challenging scenes, such as natural images with rich textures and complex backgrounds. These issues are further exacerbated when using latent diffusion models (rombach2022high), where the inclusion of the encoder-decoder architecture introduces additional challenges during the reverse process due to the nonlinearities introduced. To mitigate these limitations, further inductive biases are needed during inference, (rout2023solving).
Measurement Without REPA With REPA Ground Truth Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionDINOv2 EncoderRefer to caption Approximate reconstructed imageRepresentation AlignmentObjectiveSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockMeasurement Matching ObjectiveMLP+
Figure 1: Overview of our proposed framework. Left: Box inpainting (top row) and Gaussian deblurring (bottom row) results, where adding Repa (third column) improves the perceptual quality of reconstructed images compared to the baseline without Repa (second column). Right: Alignment mechanism between diffusion-model features and pretrained DINOv2 embeddings, combined with measurement matching. An approximate reconstructed image is provided as input to the DINOv2 encoder for computing the alignment signal.
Recent studies have shown that diffusion models learn semantic features in their hidden states, (li2023your). The more expressive these features are the better the diffusion model performs of on the generative task, (xiang2023). Building on this insight, the seminal work of yu2024repa introduced a regularizer that aligns the internal representations of the diffusion model with those of a pretrained visual encoder, (oquab2023dinov2). This framework, termed representation alignment, was shown to act as a strong semantic constraint leading to higher-fidelity generations and significantly faster convergence.
The success of Repa has attracted considerable interest in this direction, (wang2025repa; tian2025u; yao2025reconstruction; leng2025repa; wang2025learning). In particular, (wang2025repa) observed that while Repa enhances early training dynamics, its benefits diminish over time and can even degrade performance in later iterations. To address this issue, they introduced an early-stopping strategy which mitigated the late-stage degradation and further improved overall convergence speed. Other works have explored applying alignment losses directly to VAE latent spaces (yao2025reconstruction; xu2025exploring), finding that diffusion models trained on these regularized spaces converge faster. Despite these advances, most efforts focus on improving training efficiency, whereas relatively little attention has been paid to leveraging representation alignment during inference, particularly for solving inverse problems. This motivates our central question:
Can we apply representation alignment ideas to benefit existing algorithms for solving inverse problems using pretrained diffusion and flow-based models?
Contributions. Our main contributions are as follows:
We introduce a general framework for solving inverse problems with diffusion and flow-based models by enforcing alignment between internal diffusion representations and DINOv2 features. Despite the absence of ground truth, alignment with proxy reconstructions remains effective thanks to the robustness of DINOv2 features (see Figure 1). We provide theoretical results that demonstrate perceptual benefits of Repa regularizer showing a) its relation with a divergence measure defined on the DINOv2 feature space and and b) how it enhances convergence of the model’s internal representations toward those of the clean image. We demonstrate the versatility of our approach by integrating representational alignment into different existing state-of-the-art inverse problem solvers. Through extensive experiments on super-resolution, Gaussian deblurring, motion deblurring, and box inpainting, we show consistent improvements over prior methods across tasks and metrics, and further demonstrate that applying our regularizer to existing solvers allows them to match their original performance using significantly fewer discretization steps.
2 Related Work
2.1 Diffusion & Flow-based Models
Generative models based on denoising are a class of deep generative models that aim to learn a data distribution $p_{0}(x)$ by training on progressively corrupted versions of real samples (sohl2015deep; ho2020denoising; song2021denoising; liu2023flow). These models define a) a forward (noising) process, which gradually interpolates between the data distribution and a reference distribution $p_{1}(x)$ (typically a Gaussian), and b) a reverse (denoising) process, which is trained to invert this corruption and iteratively reconstruct data from noise. Following the formulation of stochastic interpolants (albergo2023stochastic), the time marginals of this process can be expressed as
$x_{t}=\alpha_{t}x^{*}+\sigma_{t}\varepsilon,\qquad x^{*}\sim p_{0}(x),\;\varepsilon\sim\mathcal{N}(0,I),$ (1)
where $\alpha_{t}$ and $\sigma_{t}$ correspond to decreasing and increasing function of $t$, respectively, with boundary conditions $\alpha_{0}=\sigma_{1}=1$ and $\alpha_{1}=\sigma_{0}=0$. Throughout, we use the notation $\dot{\alpha}_{t}:=\tfrac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$ and $\dot{\sigma}_{t}:=\tfrac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$ to denote time derivatives. lipman2023flow showed that there exists a probability flow ordinary differential equation (ODE) whose solution follows the same time marginals as the process in eq. (1):
$\dot{x}_{t}=v(x_{t},t),$ (2)
where $v(x_{t},t)$ denotes the true drift field that transports $p_{1}(x)$ to $p_{0}(x)$. To enable sampling from the data distribution, one has to train a neural network $v_{\theta}(x_{t},t)$ to approximate $v(x_{t},t)$. This is done by regressing the network output toward the optimal vector field using conditional flow matching (lipman2023flow; hyvarinen2005estimation), minimizing
$\mathcal{L}_{\mathrm{velocity}}(\theta):=\mathbb{E}_{x^{\ast},\varepsilon,t}\bigl[\|v_{\theta}(x_{t},t)-\dot{\alpha}_{t}x^{\ast}-\dot{\sigma}_{t}\varepsilon\|^{2}\bigr],$ (3)
Beyond the ODE view, there also exists a reverse stochastic differential equation (SDE) with the same time marginals:
$\mathrm{d}x_{t}=v(x_{t},t)\,\mathrm{d}t-g^{2}(t)\,s(x_{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}\bar{w}_{t},$ (4)
where $s(x_{t},t)=\nabla_{x_{t}}\log p_{t}(x_{t})$ is the score, $g(t)$ is the diffusion coefficient, and $\bar{w}_{t}$ is a standard Wiener process running backward in time. Notably, the score can be expressed directly in terms of the velocity field:
$s(x_{t},t)=\sigma_{t}^{-1}\cdot\frac{\alpha_{t}v(x_{t},t)-\dot{\alpha}_{t}x_{t}}{\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t}}.$ (5)
eq. (5) hints that the score parameterization used in diffusion models and the velocity parameterization used in flow-based models can lead to equivalent formulations of the sampling process. In practice, diffusion models correspond to a stochastic discretization of the probability flow ODE, while flow-based models correspond to its deterministic counterpart.
2.2 Diffusion Models with Representation Guidance
Diffusion models learn rich, discriminative features in their hidden states, which are crucial for their generative performance (xiang2023). Nevertheless, their learned representations still fall behind on downstream tasks when compared with state-of-the-art self-supervised visual encoders, (assran2023self; simeoni2025dinov3). This representation gap has been identified as a key bottleneck for improving the generative performance of diffusion models, as richer and more semantically structured representations have been shown to enhance sample fidelity (yu2024repa). To address this, a growing line of work introduces representation guidance, which conditions diffusion models on external representations from pretrained self-supervised encoders (li2024return). Typically these approaches fall into two categories: (i) conditioning the diffusion model on external representations, either learned, (li2024return), or retrieved from a database of samples, (blattmann2022retrieval; hu2023self; sheynin2022knn); and (ii) explicitly aligning the model’s internal representations with those of large pretrained encoders, thereby encouraging more semantically meaningful feature-spaces, (tian2025u; hu2023self; leng2025repa).
Interestingly, wang2025learning demonstrated that methods from both families can be unified under a common variational framework by deriving an evidence lower bound (ELBO) conditioned on the auxiliary representation. The main difference between these two families lies in how the auxiliary representations are integrated into the training process: in the first family, the representations are generated independently of the diffusion dynamics and serve solely as conditioning signals, whereas in the second family, the training process itself is explicitly regularized to align its intermediate features with the target representations. In particular, yu2024repa introduced Repa, a method that enforces feature-level alignment in a transformer-based diffusion model (ma2024sit), to improve sample quality and training efficiency by encouraging semantically meaningful representations to emerge inside the diffusion model.
Within the Repa framework, feature-level alignment is applied directly to intermediate representations of the diffusion model. In particular, let the diffusion model consist of $L$ transformer blocks, and $\ell\in\{1,\dots,L\}$ denote the layer at which features are extracted. The quantity $h_{t}\in\mathbb{R}^{N\times D_{2}}$ denotes the matrix of patch tokens obtained after the first $\ell$ transformer blocks when the model processes an input image $x$ at timestep $t$. Here, $N$ is the number of patch tokens and $D_{2}$ the embedding dimension, with $h_{t}^{[n]}\in\mathbb{R}^{D_{2}}$ representing the $n$-th patch token. For convenience, this is written as
$h_{t}=\textsc{DiffEnc}(x,t),$
where DiffEnc returns the patch-level hidden representation after the first $\ell$ blocks of the diffusion model’s transformer encoder. Repa enforces patch-wise similarity between the $D_{1}$-dimensional DINOv2 representation of the target image $x$, i.e., $f^{[n]}_{\text{DINOv2}}(x)\in\mathbb{R}^{D_{1}}$ for patches $n=1,\dots,N$, and the intermediate hidden states $h^{[n]}_{t}$, using the following loss:
[…]
where $g_{\phi}:\mathbb{R}^{D_{2}}\to\mathbb{R}^{D_{1}}$ is a learnable multi-layer perceptron (MLP) that projects hidden states into the DINOv2 embedding space. By jointly optimizing $\mathcal{L}_{\text{REPA}}$ with the standard diffusion loss, the model learns internal representations that are semantically richer and better aligned with high-level visual features, ultimately yielding higher-quality image generation and faster convergence during training.
2.3 Solving Inverse Problems With Diffusion and Flow-based Models
When solving inverse problems the goal is to recover the clean signal $x_{0}$ from noisy or degraded measurements $y$. In the flow- and diffusion-based framework, this is accomplished by replacing the unconditional score function in eq. (4) with the conditional score $\nabla_{x_{t}}\log p(x_{t}\mid y)$. Applying Bayes’ rule, the reverse-time SDE becomes
$dx_{t}=\Bigl(v(x_{t},t)-g^{2}(t)\bigl[\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p(y\mid x_{t})\bigr]\Bigr)dt+g(t)d\bar{w}_{t},$ (7)
The term $\nabla_{x_{t}}\log p_{t}(x_{t})$ is the unconditional score and can be computed using a pretrained diffusion/flow model. Intuitively, this term guides $x_{t}$ toward the learned data distribution. The term $\nabla_{x_{t}}\log p(y\mid x_{t})$ enforces consistency with the measurements. However, computing $p(y\mid x_{t})$ is generally intractable, as it requires marginalizing over all possible clean signals $x_{0}$.
As a result, practical algorithms rely on approximations of $\nabla_{x_{t}}\log p(y\mid x_{t})$. A recent survey by daras2024survey provides a comprehensive taxonomy of these methods. One of the most prominent approaches is Diffusion Posterior Sampling (DPS), (chung2023diffusion), which uses the following approximation:
$p(y\mid x_{t})=\int p(y\mid x_{0})p(x_{0}\mid x_{t})\,dx_{0}\approx p\left(y\mid\hat{x}_{0}=\mathbb{E}[x_{0}\mid x_{t}]\right),$ (8)
where $\mathbb{E}[x_{0}\mid x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$. In many practical scenarios, performing the diffusion process directly in pixel space is computationally prohibitive. Latent diffusion models (LDMs) address this limitation by operating in a compressed latent space. Let $\mathcal{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ be an encoder mapping an image $x_{0}$ to its latent representation $z_{0}=\mathcal{E}(x_{0})$, and let $\mathcal{D}$ denote the corresponding decoder such that $\hat{x}_{0}=\mathcal{D}(z_{0})\approx x_{0}$. Building on this formulation, recent work has proposed latent-space counterparts of pixel-based inverse problem solvers. For example, Diffusion Posterior Sampling (DPS) has been formulated in latent space by redefining the measurement likelihood as
$p(y\mid z_{t})\approx p\!\left(y\,\middle|\,\hat{x}_{0}=\mathcal{D}\!\big(\mathbb{E}[z_{0}\mid z_{t}]\big)\right).$ (9)
This straightforward formulation commonly referred to as Latent DPS has been observed to introduce artifacts, largely due to the nonlinearity of the decoder. To mitigate this effect, rout2023solving proposed a regularization term encouraging the latent variables to remain close to a fixed point of the encoder–decoder mapping, while song2024solving enforced data consistency during optimization followed by a projection of the latent variables back onto the noisy manifold.
3 Proposed Approach
Given a noisy observation $y\in\mathbb{R}^{m}$ of an unknown signal $x\in\mathbb{R}^{k}$, our goal is to sample from $p_{\theta}(x_{0}\mid y)$ using a pretrained diffusion- or flow-based model.
Given the success of the Repa framework, we aim to apply it to inverse problems. However, a key challenge is that representation alignment requires access to the ground-truth signal, which is not available in this setting. To overcome this, we must choose an alternative input to the DINOv2 encoder that can stand in for the missing ground truth. In particular, we need an approximation $\hat{x}_{0}$ of the unknown $x_{0}$ that produces a proxy representation $c_{\text{proxy}}\in\mathbb{R}^{N\times D_{1}}$, where $N$ denotes the number of patch tokens and $D_{1}$ is their feature dimension. This representation should satisfy
$c_{\text{proxy}}\equiv f_{\text{DINOv2}}(\hat{x}_{0})\approx f_{\text{DINOv2}}(x_{0}).$ (10)
Then, we sample using a regularized objective that augments the standard posterior with an additional representation-guidance term:
$J_{t}(x_{t}\mid y,c_{\text{proxy}})=\underbrace{\log p(y\mid x_{t})}_{\text{likelihood}}+\underbrace{\log p_{t}(x_{t})}_{\text{prior}}+\lambda\,\underbrace{\log p(c_{\text{proxy}}\mid x_{t})}_{\text{REPA term}},$ (11)
where $\lambda>0$ controls the strength of the additional guidance term. Here, the first term enforces measurement consistency, the second term corresponds to the unconditional prior and the last term acts as an additional semantic regularizer encouraging alignment with $c_{\text{proxy}}$ in the DINOv2 feature space.
Representation Alignment term. Let us express the flow- or diffusion-based model as $u_{\theta}=G_{1}\circ(G_{2}(x_{t},t))$, and denote $h_{t}^{[n]}=G_{2}^{[n]}(x_{t},t)$ the $n$-th patch feature extracted from the diffusion model’s hidden state $h_{t}$ at timestep $t$. Following the training loss of Repa in eq. (6), we approximate the gradient of the representation likelihood term with a sum of cosine similarities between patch embeddings:
$\nabla_{x_{t}}\log p(c_{\text{proxy}}\mid x_{t})\;\approx\;\nabla_{x_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}^{[n]}(x_{t},t))\bigr),$ (12)
where $g_{\phi}$ is a projection head mapping the model’s features into the DINOv2 feature-space (Section 2.2). When incorporated into the sampling objective, the term in eq. (12) encourages the intermediate representations of the model to align with $c_{\text{proxy}}$, i.e., the DINOv2 representation of $\hat{x}_{0}$ .
On the selection of $c_{\text{proxy}}$. To address the fact that representation alignment (yu2024repa) is not directly applicable to inverse problems where the ground-truth image is unavailable at inference time, we construct $c_{\text{proxy}}$ from an approximate reconstruction. When the measurement $y$ lies in the image domain, we simply use $f_{\text{DINOv2}}(y)$. Otherwise, we form a proxy using the best denoised estimate available at timestep $t$, namely $f_{\text{DINOv2}}(\mathbb{E}[x_{0}\mid x_{t}])$. In both cases, we rely on the robustness of pretrained DINOv2 features, which as shown in our experiments remain stable under the degradations present in super-resolution and deblurring tasks. An ablation in Appendix A.4 quantifies the reliability of this proxy across tasks.
Extension to latent diffusion models. The proposed representation-alignment strategy applies naturally to latent diffusion models, where sampling is performed in a learned latent space rather than in pixel space. Let $z_{t}$ denote the latent state at timestep $t$. In this setting, the regularized sampling objective of eq. (11) becomes
$J_{t}(z_{t}\mid y,c_{\text{proxy}})=\log p(y\mid z_{t})+\log p_{t}(z_{t})+\lambda\,\log p(c_{\text{proxy}}\mid z_{t}).$
Following the same construction as in the pixel-space formulation, we approximate the gradient of the Repa alignment term by
$\nabla_{z_{t}}\log p(c_{\text{proxy}}\mid z_{t})\;\approx\;\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}^{[n]}(z_{t},t))\bigr),$
Our general algorithm for a latent-space velocity-based sampler is summarized in Algorithm 1. Step 6 is intentionally kept abstract; as discussed in Section 2.3, different inverse-problem solvers approximate this term in different ways. The proposed regularizer is compatible with standard diffusion-based reconstruction pipelines. In our experiments, we instantiate the framework using two state-of-the-art latent diffusion methods, with further implementation details provided in Appendix A.2 and Appendix A.3.
Algorithm 1 Repa-regularized Inverse Algorithm
1:Flow model $u_{\theta}=G_{1}\circ G_{2}$, measurement $y$, learning rate $\eta$, regularizer strength $\lambda$, proxy representation $c_{\text{proxy}}$
2:initialize $z_{T}\sim\mathcal{N}(0,I)$
3:for $t\in\{T,\dots,1\}$ do
4:  $v\leftarrow u_{\theta}(z_{t},t)$
5:  $\Delta t\leftarrow 1/T$
6:  $z_{t}\leftarrow z_{t}+\Delta t\cdot v$
7:  $z_{t-1}\leftarrow z_{t}+\eta\nabla_{z_{t}}\log p(y\mid z_{t})$
8:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}(z_{t},t)^{[n]})\bigr)$
9:end for
10:return $x_{0}$
4 Theoretical Results
Recent advances in generative modeling have demonstrated that objectives defined in deep feature spaces, rather than purely pixel-based discrepancies, produce reconstructions that better capture perceptual similarity (dosovitskiy2016generating; zhang2018unreasonable; fu2023dreamsim). Motivated by this observation, our goal in this section is to clarify how the Repa regularizer fits within this broader class of feature-based objectives. In particular, we present two theoretical results that characterize how Repa acts on (i) the DINOv2-space representations of reconstructed images and (ii) the hidden representations of the diffusion model itself.
First, we show that the Repa term can be viewed as an empirical surrogate for minimizing a feature-space divergence specifically, the maximum mean discrepancy (MMD) between the DINOv2 features of reconstructed images and those of the ground truth. Second, we study how Repa affects the hidden states of the diffusion model during the reverse process. Under suitable assumptions, we show that a Repa update moves the current diffusion representation toward the representation that the model would produce for the clean image. This provides a complementary view: Repa influences not only the DINOv2 representation of the reconstruction but also the internal representations that guide the sampling trajectory.
Before presenting the main results, we introduce two quantities that capture, respectively, the alignment between feature spaces and the accuracy of the proxy representations used in our analysis.
Definition 1 (Misalignment between DINOv2 and diffusion representations).
For an image $x$ and diffusion timestep $t$, we define the representation misalignment as
$\mathrm{MisREPA}(x,t):=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-g_{\phi}\!\bigl(\mathrm{DiffEnc}^{[n]}(x,t)\bigr)\bigr\|_{2}^{2},$
where $f_{\mathrm{DINOv2}}^{[n]}(x)$ and $\mathrm{DiffEnc}^{[n]}(x,t)$ denote the $n$-th patch features from the DINOv2 encoder and the diffusion model encoder, respectively. The mapping $g_{\phi}$ is the projection head introduced in Section 2.2.
Remark.
When the timestep argument is omitted, all $t$-dependent quantities are evaluated at $t=0$. In particular, $\mathrm{MisREPA}(x)\equiv\mathrm{MisREPA}(x,0)$ and $\mathrm{DiffEnc}^{[n]}(x)\equiv\mathrm{DiffEnc}^{[n]}(x,0)$.
Definition 2 (Proxy approximation error).
For a ground-truth image $x$ and its proxy approximation $\bar{x}$ (defined in Section 3), define the proxy approximation error as
$\mathrm{ApproxErr}(x,\bar{x}):=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-f_{\mathrm{DINOv2}}^{[n]}(\bar{x})\bigr\|_{2}^{2}.$
In practice, both quantities defined above are typically small. The feature misalignment $\mathrm{MisREPA}(x,t)$ is mitigated through the explicit alignment of DINOv2 and diffusion encoder representations during training (yu2024repa), while the approximation error $\mathrm{ApproxErr}(x,\bar{x})$ is reduced by constructing proxy features $\bar{x}$ that serves as reliable substitutes for the unavailable ground-truth features (see Section 3). Thanks to the robustness of DINOv2 features to common image degradations (oquab2023dinov2), the resulting approximation gap remains small across the inverse problems we study.
Feature-Space Alignment: Repa as an MMD Surrogate
Building on the definitions introduced above, we now establish a connection between the Repa regularizer and the statistical discrepancy between DINOv2 features of clean and reconstructed images. The proposition that follows formalizes this link by expressing the expected Repa score in terms of a feature-space divergence.
Proposition 1.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$ (Section 3), with $\bar{x}\sim p_{\bar{X}}$. Define the Repa alignment as
[…]
and the empirical mean DINOv2 feature embedding
$\mu_{f}(x):=\frac{1}{N}\sum_{n=1}^{N}f_{\mathrm{DINOv2}}^{[n]}(x).$
Then the expected Repa alignment satisfies
[…]
where
$\textsc{MMD}_{\mathrm{DINOv2}}\!\left(p_{X},\,p_{\hat{X}}\right)=\bigl\|\mathbb{E}_{x}[\mu_{f}(x)]-\mathbb{E}_{\hat{x}}[\mu_{f}(\hat{x})]\bigr\|_{2}^{2}$
is the maximum mean discrepancy between $p_{X}$ and $p_{\hat{X}}$ in the DINOv2 feature space.
The proof is provided in the Appendix A.1. Proposition 1 establishes a connection between the representation alignment term and the maximum mean discrepancy (MMD) (arbel2019maximum), which quantifies perceptual divergence between feature distributions. This relationship indicates that maximizing the Repa term encourages the minimization of a divergence measure captured by the MMD together with additional penalties accounting for feature misalignment between diffusion and DINOv2 representations, as well as an approximation error arising from the use of proxy rather than ground-truth features for alignment.
When these terms are sufficiently small, Proposition 1 implies that maximizing the Repa term effectively corresponds to minimizing the maximum mean discrepancy (MMD) between the DINOv2 feature distributions of reconstructed and ground-truth images. This interpretation reveals that Repa regularization offers a principled, representation-based mechanism for enforcing perceptual consistency, by guiding generated images to align with the semantic structure captured by the DINOv2 encoder.
Effect of REPA on Internal Diffusion Representations
Having characterized the feature-space effect of Repa, we now turn to its influence on the representations of the diffusion model itself. The next result formalizes how a single Repa update impacts the model’s internal representation relative to that of the clean image.
Proposition 2.
Consider an inverse problem in which an observation $y\in\mathbb{R}^{m}$ is generated from a ground-truth image $x\sim p_{X}$, and let $\bar{x}\sim p_{\bar{X}}$ be the proxy approximation used during REPA (Section 3). Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at timestep $t$ produced by Algorithm 1 given $y$, and let $z_{t}^{\mathrm{(REPA)}}$ denote the diffusion state after applying the Repa regularization update. Let $z^{*}$ denote the latent variable associated with the clean image $x$, and define
$h_{t}=G_{2}(z_{t},t),\qquad h_{t}^{\mathrm{(REPA)}}=G_{2}(z_{t}^{\mathrm{(REPA)}},t),\qquad h^{*}=G_{2}(z^{*},0),$
where $G_{2}$ denotes the feature extractor returning the patch-level representation in layer $\ell$ of the model. Assume that the projection head $g_{\phi}$ is linear and that all DINOv2 patch embeddings are $\ell_{2}$–normalized. Then there exists a regularization weight $\lambda>0$ (Algorithm 1) such that the REPA update induces an approximate contraction, i.e., there exists a constant $C_{1}\in(0,1)$ and a constant $C_{2}\geq 0$ satisfying
$\|h_{t}^{\mathrm{(REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
The proof is provided in Appendix A.1. Proposition 2 characterizes how the Repa update modifies the hidden representations of the diffusion model. The inequality shows that a single Repa step contracts the distance between the current representation $h_{t}$ and the clean-image representation $h^{*}$, up to two additive terms: the proxy error $\mathrm{ApproxErr}(x,\bar{x})$ and the encoder misalignment $\mathrm{MisREPA}(x)$. As explained, both quantities are expected to be small in practice. When this is the case, the result implies that Repa acts as a contraction on the model’s internal representations, steering them toward those associated with the clean image.
Intuitively, Proposition 1 shows that the Repa regularizer induces a reduction in the discrepancy between reconstructed and ground-truth images in the DINOv2 feature space an embedding space known to correlate well with human perceptual similarity (fu2023dreamsim). Proposition 2 complements this view by establishing that the REPA update also reduces the gap between the model’s internal hidden representation and that of the clean image.
To further illustrate these insights, Appendix A.5 compares Repa to a variant that enforces consistency directly in the DINOv2 feature space, without aligning the diffusion model’s internal states. The results show that Repa leads to higher perceptual quality, highlighting the benefit of guiding the sampling process through the model’s intermediate representations rather than relying solely on feature-space matching.
Taken together, these findings indicate that Repa promotes perceptually meaningful reconstructions by steering the model toward regions of the feature space that reflect the semantic structure of natural images. Empirical evaluations in Section 5 reinforce this interpretation across multiple solvers and inverse problems.
5 Experimental Results
In this section, we present experimental results that demonstrate the effectiveness of our alignment regularizer. We evaluated reconstruction quality using four widely adopted metrics: PSNR (peak signal-to-noise ratio), SSIM (structure similarity index) (wang2004image), LPIPS (learned perceptual image patch similarity) (zhang2018unreasonable), and FID (Fréchet Inception Distance) (heusel2017gans). All metrics are averaged over 100 images from the ImageNet $256\times 256$ validation set (deng2009imagenet).
We consider four inverse problems: super-resolution, box inpainting, Gaussian deblurring, and motion deblurring. Following the standard corruption operators used in state-of-the-art diffusion-based inverse problem methods (chung2023diffusion; zhang2018unreasonable) we adopt the same degradation models in our experiments. For super-resolution, images are downsampled by a factor of $4$. For box inpainting, we mask out a $128\times 128$ square region. Gaussian deblurring is performed using a $61\times 61$ kernel with standard deviation $3.0$. Motion blur is simulated using randomly generated kernels with intensity $0.5$. In all experiments we also add Gaussian noise with standard deviation 0.01.
5.1 Effectiveness of the Repa Regularizer
This subsection evaluates how the Repa regularizer influences reconstruction quality by applying it to two representative latent-space solvers. To this end, we instantiate Algorithm 1 with Latent DPS and ReSample (song2024solving), which differ in how the measurement-matching term is approximated. A detailed description of these methods and their integration with Repa is provided in Appendix A.2. In all experiments, we use the latent diffusion model trained with representation alignment from yu2024repa, together with its pretrained MLP that projects internal diffusion representations into the DINOv2 feature space. Following yu2024repa, the regularizer is applied to the representations extracted after the eighth transformer block.
Table 1 reports quantitative results across all four inverse problems. Incorporating REPA yields consistent improvements in perceptual metrics notably reducing LPIPS and FID for both Latent DPS and ReSample while maintaining comparable PSNR and SSIM values. These results suggest that representation alignment guides the diffusion trajectory toward semantically consistent and perceptually realistic reconstructions that better match the target image. Figure 2 presents qualitative comparisons for box inpainting and Gaussian deblurring. Reconstructions obtained with Repa are visually sharper and more faithful to the ground truth, exhibiting fewer artifacts and clearer fine structures than those produced by the unregularized methods. Overall, these results confirm that applying Repa as an inference-time regularizer delivers both quantitative and perceptual gains, validating the proposed alignment mechanism as a general improvement for diffusion-based inverse problem solvers.
Table 1: Performance comparison of Latent DPS and ReSample (with and without the proposed Repa regularizer) on $4\times$ super-resolution, box inpainting, Gaussian deblurring, and motion deblurring tasks (ImageNet validation). Gaussian noise with $\sigma=0.01$ was added in all cases.
Task Method LPIPS ↓ FID ↓ PSNR ↑ SSIM ↑
$4\times$ Super-resolution Latent DPS 0.238 123.82 26.88 0.732
Latent DPS + REPA 0.217 86.88 26.82 0.731
Resample 0.208 97.02 26.60 0.732
Resample + REPA 0.197 74.70 26.41 0.724
Box Inpainting Latent DPS 0.151 116.53 20.53 0.827
Latent DPS + REPA 0.139 88.69 20.45 0.824
Gaussian Deblurring Latent DPS 0.288 152.96 25.76 0.648
Latent DPS + REPA 0.256 102.99 25.66 0.669
Resample 0.259 115.47 26.19 0.699
Resample + REPA 0.223 89.67 25.99 0.695
Motion Deblurring Latent DPS 0.249 129.08 27.19 0.738
Latent DPS + REPA 0.225 90.23 27.01 0.735
Resample 0.210 94.95 27.27 0.736
Resample + REPA 0.192 75.11 27.18 0.738
Measurement Latent–DPS + REPA Reference
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Box inpainting
Measurement Latent–DPS + REPA Reference
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(b) Gaussian deblurring
Figure 2: Qualitative comparison of Latent DPS with and without the proposed REPA regularizer on two inverse problems: (a) box inpainting and (b) Gaussian deblurring.
Effect of REPA with Varying Discretization Steps. Figure 3 illustrates how LPIPS varies with the number of discretization steps across three inverse problems. Incorporating the REPA regularizer consistently reduces perceptual error compared to Latent DPS alone. More importantly, REPA enables comparable or better perceptual quality with substantially fewer sampling steps: for example, in super-resolution and Gaussian deblurring, REPA achieves the same LPIPS with roughly $4\times$ fewer steps, while in motion deblurring it achieves the same level with about $2\times$ fewer steps. These results highlight the efficiency gains provided by our method, reducing computational cost without sacrificing the reconstruction quality of the baseline method.
(a) Super Resolution
(b) Gaussian Debluring
(c) Motion Debluring
Figure 3: Comparison of LPIPS as a function of discretization steps for Latent DPS and Latent DPS + REPA. Using our regularizer, comparable performance is achieved with substantially fewer number of steps.
MeasurementDPSLatent DAPSReSample + REPAGround Truth
Figure 4: Qualitative comparison for motion deblurring. Each column shows the measurement, DPS, Latent DAPS, ReSample + REPA, and the ground-truth image.
Table 2: Comparison of our REPA-regularized reconstructions against state-of-the-art inverse problem solvers on four tasks ($4\times$ super-resolution, box inpainting, Gaussian deblurring, and motion deblurring) using ImageNet validation images. Gaussian noise with $\sigma=0.01$ is added in all cases. This table complements Table 1 by benchmarking REPA-enhanced models against external state-of-the-art baselines.
Task Method LPIPS ↓ FID ↓ PSNR ↑ SSIM ↑
$4\times$ Super-resolution DPS 0.244 92.13 24.42 0.681
Latent DAPS 0.252 120.57 27.13 0.744
Latent DPS + REPA 0.217 86.88 26.82 0.731
Resample + REPA 0.197 74.70 26.41 0.724
Box Inpainting DPS 0.191 97.95 19.11 0.769
Latent DPS + REPA 0.139 88.69 20.45 0.824
Gaussian Deblurring DPS 0.366 157.73 19.55 0.461
Latent DAPS 0.300 163.36 26.02 0.692
Latent DPS + REPA 0.256 102.99 25.66 0.669
Resample + REPA 0.223 89.67 25.99 0.695
Motion Deblurring DPS 0.242 89.06 24.17 0.678
Latent DAPS 0.274 135.18 27.01 0.734
Latent DPS + REPA 0.225 90.23 27.01 0.735
Resample + REPA 0.192 75.11 27.18 0.738
5.2 Comparison with other state of the art algorithms
After establishing that the REPA regularizer improves reconstruction quality, we compare our approach with other state-of-the-art algorithms. For pixel-space methods, we consider DPS (chung2023diffusion), which performs gradient-based sampling and is evaluated using the pretrained diffusion model from dhariwal2021diffusion. For latent diffusion approaches, we include Latent DAPS. Since our method relies on a model trained on ImageNet, we evaluated Latent DAPS with the standard conditional latent diffusion model trained on ImageNet from rombach2022high, to ensure a fair comparison. All aforementioned algorithms are run with their default configurations. Table 2 presents the quantitative results, where the inclusion of our regularizer consistently achieves state-of-the-art performance. Figure 4 provides qualitative examples for motion deblurring, further highlighting the advantage of our method. In particular, DPS often introduces visible artifacts, whereas Latent DAPS fails to fully correct blur and produces reconstructions lacking sharpness.
6 Conclusions
In this work, we introduced representation alignment as a principled framework for solving inverse problems with diffusion- and flow-based generative models. By aligning internal model representations with features from a pretrained DINOv2 encoder, our method provides a strong semantic prior that guides the reconstruction process toward perceptually meaningful solutions. We further offered theoretical insights into this mechanism, showing that REPA regularization corresponds to minimizing a divergence in the DINOv2 feature space and reduces the discrepancy between the model’s internal representations and those of the clean image. These results elucidate why feature-space alignment can enhance perceptual fidelity in inverse problem settings. Overall, our study demonstrates that representation alignment constitutes a flexible and effective tool for improving generative inverse problem solvers, yielding benefits at inference time and integrate seamlessly with a range of diffusion- and flow-based methods.
Appendix A Appendix
A.1 Proofs
Proposition 0.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$ (Section 3), with $\bar{x}\sim p_{\bar{X}}$. Define the REPA alignment as
[…]
and the empirical mean DINOv2 feature embedding
$\mu_{f}(x):=\frac{1}{N}\sum_{n=1}^{N}f_{\mathrm{DINOv2}}^{[n]}(x).$
Then the expected REPA alignment satisfies
[…]
where
$\textsc{MMD}_{\mathrm{DINOv2}}\!\left(p_{X},\,p_{\hat{X}}\right)=\bigl\|\mathbb{E}_{x}[\mu_{f}(x)]-\mathbb{E}_{\hat{x}}[\mu_{f}(\hat{x})]\bigr\|_{2}^{2}$
is the maximum mean discrepancy between $p_{X}$ and $p_{\hat{X}}$ in the DINOv2 feature space.
Proof.
[…]
Assume the features are $\ell_{2}$–normalized. Then for each $n$,
$\cos\!\left(f_{\text{DINOv2}}^{[n]}(\bar{x}),\,g_{\phi}\!\big(\textsc{DiffEnc}^{[n]}(\hat{x})\big)\right)=1-\tfrac{1}{2}\,\big\|f_{\text{DINOv2}}^{[n]}(\bar{x})-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big\|_{2}^{2},$
and hence
[…]
Let the empirical (mean) feature embeddings be
$\mu_{f}(x)$ $:=\frac{1}{N}\sum_{n=1}^{N}f_{\text{DINOv2}}^{[n]}(x),$ $\mu_{f}(\bar{x})$ $:=\frac{1}{N}\sum_{n=1}^{N}f_{\text{DINOv2}}^{[n]}(\bar{x}),$
[…]
From Jensen’s inequality for the squared norm :
$\big\|\mu_{f}(x)-\mu_{f}(\hat{x})\big\|_{2}^{2}\leq 2\,\big\|\mu_{f}(x)-\mu_{g}(\hat{x})\big\|_{2}^{2}+2\,\big\|\mu_{f}(\hat{x})-\mu_{g}(\hat{x})\big\|_{2}^{2}.$
Moreover,
$\big\|\mu_{f}(x)-\mu_{g}(\hat{x})\big\|_{2}^{2}=\Big\|\frac{1}{N}\sum_{n=1}^{N}\!\big(f_{\text{DINOv2}}^{[n]}(x)-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big)\Big\|_{2}^{2}\;\leq\;\frac{1}{N}\sum_{n=1}^{N}\!\big\|f_{\text{DINOv2}}^{[n]}(x)-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big\|_{2}^{2},$
and similarly
[…]
Combining the two yields
[…]
Using the cosine–squared identity for unit–norm features,
$),\frac{1}{N}\sum_{n=1}^{N}\big\|f_{\text{DINOv2}}^{[n]}(\bar{x})-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big\|_{2}^{2}=2\big(1-\text{REPA}(\bar{x},\hat{x}$$\big)$,
[…]
Rearranging gives the bound
[…]
Taking expectations over $x\sim p_{X}$, and $\hat{x}\sim p_{\hat{X}}$, and using the convexity of the squared norm,
[…]
we obtain
[…]
Proposition 1.
Consider an inverse problem in which an observation $y\in\mathbb{R}^{m}$ is generated from a ground-truth image $x\sim p_{X}$, and let $\bar{x}\sim p_{\bar{X}}$ be the proxy approximation used during REPA (Section 3). Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at timestep $t$ produced by Algorithm 1 given $y$, and let $z_{t}^{\mathrm{(REPA)}}$ denote the diffusion state after applying the Repa regularization update. Let $z^{*}$ denote the latent variable associated with the clean image $x$, and define
$h_{t}=G_{2}(z_{t},t),\qquad h_{t}^{\mathrm{(REPA)}}=G_{2}(z_{t}^{\mathrm{(REPA)}},t),\qquad h^{*}=G_{2}(z^{*},0),$
where $G_{2}$ denotes the feature extractor returning the patch-level representation in layer $\ell$ of the model. Assume that the projection head $g_{\phi}$ is linear and that all DINOv2 patch embeddings are $\ell_{2}$–normalized. Then there exists a regularization weight $\lambda>0$ (Algorithm 1) such that the REPA update induces an approximate contraction, i.e., there exists a constant $C_{1}\in(0,1)$ and a constant $C_{2}\geq 0$ satisfying
$\|h_{t}^{\mathrm{(REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
Proof. We first rewrite all quantities in vectorized form. Let $\mathrm{vec}(\cdot)$ denote the operator that stacks the rows of a matrix into a single vector. Thus, for any matrix $H\in\mathbb{R}^{k_{1}\times k_{2}}$ we define
$\widetilde{H}:=\mathrm{vec}(H)\in\mathbb{R}^{k_{1}k_{2}}.$
Applying this to our setting, for the DINOv2 features $f_{\mathrm{DINOv2}}(x)\in\mathbb{R}^{N\times D_{1}}$ we write
$\widetilde{f}(x):=\mathrm{vec}\!\left(f_{\mathrm{DINOv2}}(x)\right)\in\mathbb{R}^{ND_{1}}.$
For the diffusion-model feature extractor $G_{2}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{N\times D_{2}}$ we define its vectorized form
$\widetilde{G}_{2}(z_{t}):=\mathrm{vec}\!\left(G_{2}(z_{t},t)\right)\in\mathbb{R}^{ND_{2}}.$
Accordingly, we set
$\widetilde{h}_{t}:=\widetilde{G}_{2}(z_{t}),\qquad\widetilde{h}_{t}^{(\mathrm{repa})}:=\widetilde{G}_{2}(z_{t}^{(\mathrm{repa})}),\qquad\widetilde{h}^{*}:=\widetilde{G}_{2}(z^{*}).$
The linear projection head $g_{\phi}(h)=\Phi h$ with $\Phi\in\mathbb{R}^{D_{1}\times D_{2}}$ is lifted patch-wise to the block-diagonal operator
$\widetilde{\Phi}:=I_{N}\otimes\Phi\in\mathbb{R}^{ND_{1}\times ND_{2}},$
where $\otimes$ denotes the Kronecker product and $I_{N}$ is the $N\times N$ identity matrix. Since DINOv2 features are $\ell_{2}$–normalized, maximizing the REPA cosine similarity is equivalent to minimizing a squared $\ell_{2}$ distance in the feature space. Thus, the REPA update step applied to $z_{t}$ can be written as
$z_{t}^{(\mathrm{repa})}=z_{t}-\lambda\,\nabla_{z_{t}}\Bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\Bigr\|_{2}^{2}.$
The corresponding diffusion representation is
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{G}_{2}\!\left(z_{t}-\lambda\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}\right).$
Applying a first–order Taylor expansion of $\widetilde{G}_{2}$ around $z_{t}$ for sufficiently small $\lambda$ yields
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{h}_{t}-\lambda J_{\widetilde{G}_{2}}(z_{t})\,\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}+O(\lambda^{2}),$ (16)
where $J_{\widetilde{G}_{2}}(z_{t})\in\mathbb{R}^{ND_{2}\times d}$ is the Jacobian of $\widetilde{G}_{2}$ at $z_{t}$.
The inner gradient is, by the chain rule,
$\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\Phi\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}=-2\,J_{\widetilde{G}_{2}}(z_{t})^{\!\top}\,\widetilde{\Phi}^{\top}\,\Big(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\Big).$
Substituting this into equation 16 gives
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{h}_{t}+2\lambda\,J_{\widetilde{G}_{2}}(z_{t})\,J_{\widetilde{G}_{2}}(z_{t})^{\!\top}\,\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
Let
$A_{t}:=J_{\widetilde{G}_{2}}(z_{t})J_{\widetilde{G}_{2}}(z_{t})^{\top}\;\in\;\mathbb{R}^{ND_{2}\times ND_{2}}.$
Subtracting $\widetilde{h}^{*}$ on both sides yields
$\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}$ $=(\widetilde{h}_{t}-\widetilde{h}^{*})+2\lambda\,A_{t}\,\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
We now add and subtract $\widetilde{f}(x)$ and $\widetilde{\Phi}\widetilde{h}^{*}$ inside the parentheses:
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{f}(x)\bigr)$
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{\Phi}\widetilde{h}^{*}-\widetilde{\Phi}\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
Grouping terms gives
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{f}(x)\bigr)$
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\bigr)+O(\lambda^{2}).$
Taking Euclidean norms and using the triangle inequality yields
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}$ $\leq\|I-2\lambda A_{t}\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}$
$\quad+2\lambda\|A_{t}\widetilde{\Phi}^{\top}\|_{2}\Bigl(\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}+\|\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\|_{2}\Bigr)+O(\lambda^{2}).$
By the definition of $\mathrm{ApproxErr}$,
$\mathrm{ApproxErr}(x,\bar{x})=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-f_{\mathrm{DINOv2}}^{[n]}(\bar{x})\bigr\|_{2}^{2},$
and in vectorized form this implies
$\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}^{2}=\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(\bar{x})-f_{\mathrm{DINOv2}}^{[n]}(x)\bigr\|_{2}^{2}=N\,\mathrm{ApproxErr}(x,\bar{x}),$
so
$\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}=\sqrt{N}\,\sqrt{\mathrm{ApproxErr}(x,\bar{x})}.$
Similarly, from the definition of $\mathrm{MisREPA}$ we obtain
$\|\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\|_{2}=\sqrt{N}\,\sqrt{\mathrm{MisREPA}(x)}.$
Substituting into the previous inequality gives
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}\leq\|I-2\lambda A_{t}\Phi^{\top}\Phi\|_{2}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}+2\lambda\sqrt{N}\,\|A_{t}\Phi^{\top}\|_{2}\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right)+O(\lambda^{2}).$
Since $A_{t}=J_{\widetilde{G}_{2}}(z_{t})J_{\widetilde{G}_{2}}(z_{t})^{\top}$ and $\widetilde{\Phi}^{\top}\widetilde{\Phi}$ are positive semidefinite,
$\|A_{t}\|_{2}=\sigma_{\max}(J_{\widetilde{G}_{2}}(z_{t}))^{2},\qquad\|\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}=\sigma_{\max}(\widetilde{\Phi})^{2},$
where $\sigma_{\max}(\cdot)$ denotes the largest singular value. Since $\widetilde{\Phi}=I_{N}\otimes\Phi$, its singular values are exactly those of $\Phi$ (each repeated $N$ times), hence $\|\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}=\sigma_{\max}(\Phi)^{2}$.
Thus, choosing
$0<\lambda<\frac{1}{\sigma_{\max}(J_{\widetilde{G}_{2}}(z_{t}))^{2}\,\sigma_{\max}(\Phi)^{2}}$
ensures
$\|I-2\lambda A_{t}\Phi^{\top}\Phi\|_{2}\leq C_{1}\qquad\text{for some }0<C_{1}<1.$
Letting
$C_{2}:=2\lambda\,\sqrt{N}\,\|A_{t}\Phi^{\top}\|_{2},$
and absorbing the $O(\lambda^{2})$ term into the constants for sufficiently small $\lambda$, we obtain the vectorized bound
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}\;\leq\;C_{1}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
Finally, since for any $H\in\mathbb{R}^{k_{1}\times k_{2}}$ we have
$\|H\|_{F}=\|\mathrm{vec}(H)\|_{2},$
the same inequality holds in matrix form with Frobenius norms:
$\|h_{t}^{(\mathrm{REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right),$
which is exactly the claim of the proposition.
A.2 Details About Latent DPS + REPA Implementation
We implement Latent DPS + REPA following the procedure in Algorithm 2. A key design choice is the learning rate schedule $\eta_{t}$, for which we consider two adaptive parameterizations:
$\eta(t)=\left\{\begin{aligned} &\frac{\kappa}{\|y-A\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\|_{2}}&&\text{(Adaptive inverse norm)}\\[12.0pt] &\frac{\kappa}{\max\!\left(\tfrac{t}{1-t},\,1\right)}&&\text{(Adaptive SNR-based learning rate)}\end{aligned}\right$.
where $\kappa$ is a tunable scaling factor. Table 3 compares these schedules for $4\times$ super-resolution on ImageNet, showing that the SNR-based strategy consistently achieves better perceptual performance (lower LPIPS) and slightly higher PSNR. For each schedule, $\kappa$ and $\lambda$ are tuned on a small validation set. We use the SNR-based learning rate for all main experiments unless otherwise.
Table 3: Comparison of learning rate strategies for $4\times$ super-resolution on ImageNet.
(a) Latent DPS
LR Type $\kappa$ PSNR ↑ LPIPS ↓
Inverse norm 7.75 26.09 0.274
SNR-based 2.25 26.82 0.238
(b) Latent DPS + REPA
LR Type $\lambda$ $\kappa$ PSNR ↑ LPIPS ↓
Inv norm 0.005 5 25.90 0.235
SNR-based 0.01 2 26.82 0.217
Algorithm 2 Latent DPS + REPA Algorithm
1:flow model $u_{\theta}$ = $G_{1}\circ G_{2}$, measurement $y$, pretrained encoder $f$
2:initialize $x_{T}\sim\mathcal{N}(0,I)$
3:for $t\in\{T,\dots,0\}$ do
4:  $v\leftarrow u_{\theta}(z_{t},t)$
5:  $\hat{z}_{0}\leftarrow\mathbb{E}[z_{0}\mid z_{t}]$
6:  $\Delta t\leftarrow 1/T$
7:  $z_{t-1}\leftarrow z_{t}+\Delta t\cdot v$
8:  $z_{t-1}\leftarrow z_{t-1}-\eta\nabla_{z_{t}}\|y-\mathcal{A}(\mathcal{D}(\hat{z}_{0}))\|_{2}^{2}$
9:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos((f_{\text{DINOv2}}^{[n]}(y),g_{\phi}(G_{2}(z_{t},t)^{[n]}))$
10:end for
11:return $x_{0}$
A.3 Details About Resample + REPA Implementation
In Table 3, we present an adaptation of the algorithm proposed by song2024solving to the flow-based setting, augmented with the REPA regularizer as described in the methodology section. We note that the original ReSample algorithm employs a three-stage procedure for enforcing data consistency: (i) gradient steps in latent space, similar to Latent DPS; (ii) pixel-space optimization, which is computationally efficient and captures high-level semantics but often leads to blurrier reconstructions; and (iii) latent space optimization as outlined to line 7 of Algorithm 3. In contrast, our adaptation omits the pixel-space stage, as our focus is on maximizing perceptual quality. We found that this modification together with the inclusion of the REPA reguralizer yields sharper and more visually convincing results, while still benefiting from the refinement effect of the final latent-space consistency updates.
Importantly, the resulting algorithm remains consistent with the central principle of ReSample—that initialization strongly impacts the effectiveness of data-consistency updates. In our case, REPA guides the initialization toward reconstructions with stronger perceptual quality, which the subsequent consistency updates then refine further, improving fidelity and sharpness.
Algorithm 3 Latent DPS + REPA Algorithm
1:flow model $u_{\theta}=G_{1}\circ G_{2}$, measurement $y$, pretrained encoder $f$, resample steps $C$, parameter $\gamma$
2:initialize $z_{T}\sim\mathcal{N}(0,I)$
3:for $t\in\{T,\dots,0\}$ do
4:  $v\leftarrow u_{\theta}(z_{t},t)$
5:  $\hat{z}_{0}\leftarrow\mathbb{E}[z_{0}\mid z_{t}]$
6:  $\Delta t\leftarrow 1/T$
7:  $z_{t-1}\leftarrow z_{t}+\Delta t\cdot v$
8:  if $t\in C$ then
9:   $\tilde{z}_{0}(y)\leftarrow\arg\min_{z}\tfrac{1}{2}\|y-\mathcal{A}(\mathcal{D}(z))\|_{2}^{2}$ $\triangleright$ initialize at $\hat{z}_{0}$
10:   $z_{t-1}\leftarrow\textsc{StochasticResample}\!\left(\tilde{z}_{0}(y),\,z_{t-1},\,\gamma\right)$
11:  end if
12:  $z_{t-1}\leftarrow z_{t-1}-\eta\nabla_{z_{t}}\|y-\mathcal{A}(\mathcal{D}(\hat{z}_{0}))\|_{2}^{2}$
13:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\big((f_{\text{DINO}}^{[n]}(y),\,g_{\phi}(G_{2}(z_{t},t)^{[n]})\big)$
14:end for
[…]
16:return $x_{0}$
A.4 Robustness of DINOv2 Representations to Corruptions
To evaluate the robustness of DINOv2 representations under various corruptions, we conduct experiments on a fixed set of 100 images. For each image, we compute the average patch similarity between its ground truth representation and that of its corrupted version. We assess this similarity across different corruption types specifically super-resolution and Gaussian deblurring at varying levels of severity. We report the average similarity across the selected subset of 100 images.
Figure 5 presents how DINOv2 representation similarity changes as the corruption severity increases for the two tasks. Figure 6 further visualizes how the pixel-space appearance of a single image changes across corruption levels. Interestingly, we observe that DINOv2 representations remain significantly more robust to both super-resolution and Gaussian deblurring. Despite substantial visual degradation in pixel space, DINOv2 maintains a strong representation alignment with the original image.
(a) Super-Resolution
(b) Gaussian Deblurring
Figure 5: Similarity of representations under increasing levels of corruption.
(a) Ground Truth
(b) Super Resolution ×4
(c) Super Resolution ×8
(d) Super Resolution ×16
(e) Ground Truth
(f) Gaussian deblur $\sigma=3$
(g) Gaussian deblur $\sigma=6$
(h) Gaussian deblur $\sigma=9$
Figure 6: Visual comparison of a corrupted image used in the similarity experiments for different corruption types and severity levels.
A.5 Enforcing Feature-Level DINOv2 Consistency without Representation Alignment
In this ablation study, we compare REPA against a variant that enforces consistency directly in the DINOv2 feature-space, without involving the diffusion model representations. More specifically, instead of using the representation alignment loss in line 8 of Algorithm 2, we replace it with the following update step:
$z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\left(f_{\text{DINOv2}}^{[n]}(y),\,f_{\text{DINOv2}}^{[n]}\!\big(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\big)\right),$ (17)
While both objectives influence the sampling trajectory, the feature-only variant derives its guidance from the DINOv2 features of the decoded image. In contrast, REPA aligns the diffusion model’s internal representations with DINOv2 features directly at each timestep.
As shown in Table 4, the direct feature-space alignment yields consistently lower perceptual quality than the proposed representation-based formulation. for both experiments we have chosen the best hyperparameters using a small grid search as in A.2 to ensure fairness in comparison. Figure 7 further illustrates this difference: REPA reconstructions are sharper and more faithful to the ground truth for both super-resolution and Gaussian deblurring. These results demonstrate that aligning the diffusion model’s internal representations provides a stronger and more semantically guided inductive bias than feature-space matching alone.
MeasurementFeature-Space AlignmentREPA (Ours)Ground Truth
Figure 7: Qualitative comparison between feature-space and representation-space alignment. Results are shown for super-resolution (top row) and Gaussian deblurring.
Table 4: Quantitative comparison between REPA and direct feature-space alignment. LPIPS and FID results for super-resolution and Gaussian deblurring on the ImageNet validation set. Qualitative examples for the same tasks are shown in Figure 7.
Super-resolution ($\times 4$) Gaussian deblurring
Method LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$
Feature-space alignment 0.227 95.69 0.276 128.29
Representation alignment (proposed) 0.217 86.88 0.256 102.99
A.6 Additional Qualitative Results
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Figure 8: Qualitative comparison for Box Inpainting. Each row shows: Measurement, pixel-space method DPS, latent method DPS, its REPA-enhanced version, and the Ground Truth.
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Figure 9: Qualitative comparison for $4\times$ super-resolution. Each row shows: Measurement, a baseline method (DPS or DAPS), a latent solver (Latent DPS or ReSample), its REPA-enhanced version, and the Ground Truth.
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Figure 10: Qualitative comparison for Gaussian Deblurring. Each row shows: Measurement, a baseline method (DPS or DAPS), a latent solver (Latent DPS or ReSample), its REPA-enhanced version, and the Ground Truth.
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Figure 11: Qualitative comparison for Motion Deblurring. Each row shows: Measurement, a baseline method (DPS or DAPS), a latent solver (Latent DPS or ReSample), its REPA-enhanced version, and the Ground Truth.